1) 1/2q-5>3Q
This is what i have so far and am stuck.
1q-5>3q
2) 13x-5=x+1
3)Not sure if the answer is right. But I did tried
3/4q+11=4q+37
3q+44=16q+148
16q+148=3q+44
148 +60
16q = 3q+60
-3 -3
13 = 60
Divdie both sides by 13 answer is -4
4) 2/3y-7=5y+19
3y-21=15y+57
15y+57=3y-21
57 +57
15y = 3y-36
-3 -3
12y = 36
divide by 12 on both sides Answer 3
5)5z-3>3z+1

Answer by stanbon(75887)

1) (1/2)q-5>3q
Multiply both sides by 2 to get:
q - 10 > 6q
---
5q< -10
---
q < -2
===================
2) 13x-5=x+1
12x = 6
x = 1/2
===================

(3/4)q+11 = 4q+37
Multiply thru by 4 to get:
3q + 44 = 16q + 148
13q = -104
q = -8
-------------------------
4) (2/3)y-7=5y+19
Multiply thru by 3 to get:
2y - 14 = 10y + 38
8y = -52
y = -13/2
---------------------------
5)5z-3>3z+1
2z > 4
z > 2
--------------
Cheers,
Stan H.